OpenStudy (anonymous):

To study the convergence of the series: ∑_(n≥1)▒〖(-1)〗^(n ) 〖(lnn)〗^2/√n

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

Can you try using Equation Editor

OpenStudy (zarkon):

\[\sum_{n=1}^{\infty}(-1)^n\frac{(\ln(n))^2}{\sqrt{n}}\] this?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

it is conditionally convergent :)

OpenStudy (zarkon):

Apply the alternating series test

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Apply the alternating series test.no

OpenStudy (zarkon):

why not

OpenStudy (anonymous):

To study the convergence of the series:(séries divergence or convergence)

OpenStudy (zarkon):

I'm not totally following what you are saying. If you apply the alternating series test you will see that the above series is convergent (though just conditionally)

OpenStudy (anonymous):

may be

OpenStudy (anonymous):

Why We do not apply the theory leibuitz

OpenStudy (anonymous):

How is the solution?please

OpenStudy (zarkon):

which theory would that be

OpenStudy (anonymous):

isn't liebniz the same as alternate series or

OpenStudy (zarkon):

it was discovered by Leibniz, but Leibniz discovered many things :)